OpenStudy (mendicant_bias):

I'm trying to figure out a simple algebraic rearrangement, but I just can't get my head around it. One moment.

OpenStudy (mendicant_bias):

\[\frac{ 1 }{ 1 + \frac{ 3 }{ \sqrt{13} }} = ?\] I can't find the original solution (It's lost forever), but it had two terms in the numerator being added, involved a 13, a sqrt(13), and a 4.

OpenStudy (mendicant_bias):

Pardon, that is supposed to be a minus sign in the denominator

OpenStudy (whpalmer4):

\[\frac{1}{1-\frac{3}{\sqrt{13}}}\]Is that it?

OpenStudy (whpalmer4):

I would start by multiplying by the fraction \[\frac{1+\frac{3}{\sqrt{13}}}{1+\frac{3}{\sqrt{13}}}\]

OpenStudy (whpalmer4):

that will give you \[\frac{1+\frac{3}{\sqrt{13}}}{1-\frac{9}{13}}\]

OpenStudy (mendicant_bias):

pI can't respond fully at the moment, but I see it; thanks. \derp

OpenStudy (radar):